# Bar Evolution in Edge-on Galaxies: A Demographic Study of Boxy/Peanut Bulges

### v1

@@ 7.1 Mass and size distributions of BP/X bulges
First, we investigate the differences in properties like mass, disk radius, disk height and axial ratio among the different classes we have identified in the whole sample. Fig. 4, we plot the distributions of mass, axis ratio, disk height, and disk radius for for samples in this study. The control, both-sided, one-sided, and strong both-sided are respectively shown with blue, green, red, and purple colors. To better understand the distribution, we fitted each dataset using the ’log normal’ distribution. For fitting, we used the Python package scipy .777scipy
The distribution of mass (4a) shows that the mass distribution for the samples without any signature of bulge formation (control sample) lies in a range of . But for the other 3 samples, i.e. the ones showing signatures of bulge formation goes up to nearly . This is expected because bulge formation occurs in galaxies having a bar and high mass galaxies are more likely to have bars than the lower mass galaxies (Erwin (2018)).
We have defined axial ratio as the ratio of disk height to disk radius. For the galaxies having signature of bulge formation, we expect the axial ratio to be higher than that of the control sample as the disk height is higher for them. For our lognorm fit, we obtained the mean value of axial ratio of strong both-sided galaxies is higher among all samples (4b).
For all our further analysis described in section 7.4 and 7.5, we have used only the galaxies having mass, disk radius and disk height less than , 14kpc and 5kpc respectively. By using these limits, we were able to remove the relatively higher values in the data so that we could study the remaining data in a more detailed manner. These limits do not affect the correctness of the analysis because this includes nearly 98.8% of the whole galaxies. Category-wise, we have more than 99% of control sample and one-sided sample , nearly 98.3% of both-sided and 97% of strong both-sided inside this limit.
@@ 7.2 BP/X fraction
Lütticke et al. (2000) presented the first detailed statistics of BP/X bulges using a large sample of edge-on galaxies from Third Reference Catalogue of Bright Galaxies (RC3: de Vaucouleurs et al., 1991). They found BP/X bulges in their sample of 734 classifiable bulges. Yoshino & Yamauchi (2015) decomposed about 1700 SDSS DR7 disc galaxies, and identified BP/X structure. In their clean sample of 1329 galaxies in band, they reported BP/X structures. Since Yoshino & Yamauchi (2015) used SDSS sample, we cross-matched our sample with the their catalog of 1312 edge-on galaxies in the -band, and identified 439 galaxies in common. The primary reason for the discrepancy in sample sizes lies in the differing selection criteria.Yoshino & Yamauchi (2015) selected galaxies with a minor-to-major isophotal axis ratio in the -band less than 0.25, whereas our sample was selected based on volunteer votes. When we compare the values from their model fitting for both the common and complete samples, we find that the common sample predominantly consists of galaxies with , while complete sample show median , indicating that a large fraction of their sample comprises moderately inclined systems. Nonetheless, out of these common galaxies they identified 123 total BP/X structures (7 strong, 40 standard, and 76 weak in their definitions) and we identified 150 total BP/X structures (65 strong both-sided and 85 both-sided). Recently, a similar fraction of BP/X bulges is reported in -band images. Marchuk et al. (2022) trained Artificial Neural Network (ANN) to automatically identify BP/X structures in large sample of 13048 edge-on galaxies obtained from DESI Legacy Imaging Survey (Dey et al., 2019). They found about 1925 () galaxies with BP/X structures.
As described previously in Section 4, we started with 6653 edge-on galaxies and classified them into three classes of BP/X bulges: Both-sided, Strong both-sided, and One-sided. We identify 1675 () both-sided bulges, and 1108 () one-sided bulges. Among both-sided bulges 545 () are strong both-sided with prominent X-structure, and remaining 1130 () show faint X-structure. When combined together, both-sided and one-sided make (2783 galaxies) of the total sample. Though our base sample is nearly half of that studied in Marchuk et al. (2022), we have visually verified the presence of BP/X structure in each galaxy by manually varying intensity of the residual images to avoid missing any faint structure. Therefore, our analysis presents largest and complete sample of BP/X structures till date with detailed classification.
@@ 7.3 Relationship of BP/X bulge with the stellar mass
In Fig. 4(a), we have plotted the stellar mass distribution of both-sided (gold), strong both-sided (green), one-sided (black), and control sample (cyan) in this study. We found that the distribution of control galaxies peaks at lower stellar mass than the galaxies with BP/X bulges. The median of the control sample lies at , whereas the sample with both-sided, strong both-sided, and one-sided BP/X bulges have median at , respectively. Our results are consistent with the previous findings of Erwin & Debattista (2017); Kruk et al. (2019); Marchuk et al. (2022), suggesting massive galaxies most likely host the BP/X bulges.
Further, we measure the BP/X fraction in our sample of edge-on galaxies. Fig. 5 shows the fraction of both-sided (green), strong both-sided (orange), one-sided (blue) bulges with the stellar mass of galaxies. The vertical bars in respective curve represent statistical uncertainty in measurement defined as , where being fraction of a bulge class at any stellar mass bin and being total number of galaxies in that bin. Similar to the previous studies, we notice increasing BP/X bulge fraction with increasing stellar mass of galaxies. In the stellar mass range shown here, the fraction of one-sided BP/X bulges is always higher than the strong both-sided bulges at any given stellar mass. Additionally, the fraction of both-sided BP/X bulges, which also includes strong both-sided bulges, is always the highest for any stellar mass. In the shown stellar mass range, we can see an increase in strong both-sided, one-sided and both-sided bulge fraction, respectively, from 0, 0, and 12 to 18, 24, and 38 percent.
@@ 7.4 Dependence of BP/X bulge on stellar surface density
Fig. 6 clearly shows that the fraction of any type of BP/X structure increases with the increasing average surface density of galaxies. The low surface density disks () show nearly constant fraction of both-sided and strong both-sided BP/X bulges. Below this surface density limit, the fraction of strong both-sided structures is less than , while both-sided makes about , and the contribution of one-side structures rises from zero to . For high surface density disks (), the fraction of BP/X structure hosting disks rises by about in all three groups. This analysis shows that the high surface density disks are more likely to host BP/X structures as compare to the low surface density disks.
@@ 8 Summary
We visually classified 6653 edge-on galaxies into 1675 () both-sided bulges, and 1108 () one-sided bulges. Among both-sided bulges 545 () are strong both-sided with prominent X-structure, and remaining 1130 () show faint X-structure. When combined together, both-sided and one-sided make (2783 galaxies) of the total sample. Our sample provides the largest catalog of BP/X bulges to date, which includes a category of one-sided BP/X bulges. The catalog of one-sided structures is one of its kind sample, which likely contains samples of ongoing buckling events in edge-on galaxies.
The fraction of one-sided BP/X bulges is always higher than the strong both-sided bulges at any given stellar mass. The fraction of both-sided BP/X bulges is always the highest for any stellar mass. At , the fraction of strong both-sided, one-sided and both-sided bulge is 18, 24, and 38 percent, respectively.
@@
This study provides a classification of BP/X bulges in a large sample of edge-on galaxies, along with their statistical properties and evolutionary trends. We demonstrate the effect of bar formation and evolution on the mass–size and mass–density scaling relations. The analysis uses single-component fitting for size measurements and empirical relations for dust correction and mass estimation, which introduce significant scatter in the scaling relations. Future studies employing detailed 2D decomposition for scale size measurements and SED fitting for stellar mass, star formation rate, dust content, and star formation history will be instrumental in the era of large-scale surveys, e.g., Euclid, LSST, 4MOST, and Roman. Additionally, our sample can be used to train AI models for studying large datasets.
@@ Acknowledgements.
- D´ıaz-Garc´ıa et al. (2020) Díaz-García, S., Moyano, F. D., Comerón, S., et al. 2020, A&A, 644, A38
- D´ıaz-Garc´ıa et al. (2016) Díaz-García, S., Salo, H., Laurikainen, E., & Herrera-Endoqui, M. 2016, Astronomy & Astrophysics, 587, A160
@@
- Gonzalez et al. (2017) Gonzalez, O. A., Debattista, V. P., Ness, M., Erwin, P., & Gadotti, D. A. 2017, Monthly Notices of the Royal Astronomical Society: Letters, 466, L93
@@
- Marchuk et al. (2022) Marchuk, A. A., Smirnov, A. A., Sotnikova, N. Y., et al. 2022, Monthly Notices of the Royal Astronomical Society, 512, 1371

### v2

@@ 6 Sample characteristics
Since the identification of BP/X structures is affected by the inclination of galaxies, it is worth discussing the inclination distribution of our sample. We measure the inclination angle using the axis ratio obtained from SDSS data, applying the relation (Holmberg, 1958), where is the intrinsic axial ratio. Following Tully & Fisher (1977), we adopt for this analysis. We find a median inclination of degree, with a – range of – degrees. For an ideally edge-on galaxy, the inclination angle is degree, indicating that our sample largely consists of nearly edge-on galaxies.
@@ 7.1 Mass and size distributions of BP/X bulges
First, we investigate the differences in properties like mass, disk radius, disk height and axial ratio among the different classes we have identified in the whole sample. In Fig. 4, we plot the distributions of mass, axis ratio, disk height, and disk radius for the four samples in this study. The control, both-sided, one-sided, and strong both-sided are respectively shown with blue, green, red, and purple colors. To better understand the distribution, we fitted each dataset using the ‘log normal’ distribution. For fitting, we used the Python package Scipy.
The distribution of masses (4a) shows that the mass distribution for the sample without any signature of bulge formation (control sample) peaks at . But for the other 3 samples, i.e. the ones showing signatures of bulge formation peaks in the range of . This is expected because bulge formation occurs in galaxies having a bar and high mass galaxies are more likely to host bars than the lower mass galaxies (Erwin (2018)).
We have defined axial ratio as the ratio of disk scale height to disk scale radius. For the galaxies having signature of bulge formation, we expect the axial ratio to be higher than that of the control sample as bars thicken during BP/X bulge formation. From our ‘lognorm fit’, we obtained the mean value of axial ratio of strong both-sided galaxies higher among all samples (4b), though not significantly different. This similarity in the axial ratio is also evident from the scale height (4c) and scale radius distributions (4d).
For further analysis described in Section 7.4 and 7.5, we have used only the galaxies having stellar mass, disk radius and disk height less than , 14 kpc, and 5 kpc, respectively. By using these limits, we remove relatively higher value outliers in the data. These limits do not affect the correctness of the analysis because the sample includes nearly 98.3% of galaxies. Category-wise, we have nearly 99% of control sample and one-sided sample, nearly 97.4% of both-sided and 95.4% of strong both-sided within this limit.
@@ 7.2 BP/X fraction
Lütticke et al. (2000) presented the first detailed statistics of BP/X bulges using a large sample of edge-on galaxies from Third Reference Catalog of Bright Galaxies (RC3: de Vaucouleurs et al., 1991). They found BP/X bulges in their sample of 734 classifiable bulges. Yoshino & Yamauchi (2015) decomposed about 1700 SDSS DR7 disc galaxies, and identified BP/X structure. In their clean sample of 1329 galaxies in band, they reported BP/X structures. Since Yoshino & Yamauchi (2015) used SDSS sample, we cross-matched our sample with the their catalog of 1312 edge-on galaxies in the -band, and identified 455 galaxies in common. The primary reason for the discrepancy in sample sizes lies in the differing selection criteria.Yoshino & Yamauchi (2015) selected galaxies with a minor-to-major isophotal axis ratio in the -band less than 0.25, whereas our sample was selected based on volunteer votes. When we compare the values from their model fitting for both the common and complete samples, we find that the common sample predominantly consists of galaxies with , while complete sample show median , indicating that a large fraction of their sample comprises moderately inclined systems. Nonetheless, out of these common galaxies they identified 129 total BP/X structures (7 strong, 42 standard, and 80 weak in their definitions) and we identified 165 total BP/X structures (68 strong both-sided and 97 both-sided). Recently, a similar fraction of BP/X bulges is reported in -band images. Marchuk et al. (2022) trained Artificial Neural Network (ANN) to automatically identify BP/X structures in large sample of 13048 edge-on galaxies obtained from DESI Legacy Imaging Survey (Dey et al., 2019). They found about 1925 () galaxies with BP/X structures.
As described previously in Section 4, we started with 6653 edge-on galaxies and classified them into three classes of BP/X bulges: Both-sided, Strong both-sided, and One-sided. We identify 1673 () both-sided bulges, and 1112 () one-sided bulges. Among both-sided bulges 504 () are strong both-sided with prominent X-structure, and remaining 1169 () show faint X-structure. When combined together, both-sided and one-sided make (2785 galaxies) of the total sample. Though our base sample is nearly half of that studied in Marchuk et al. (2022), we have visually verified the presence of BP/X structure in each galaxy by manually varying intensity of the residual images to avoid missing any faint structure. Therefore, our analysis presents largest and complete sample of BP/X structures till date with detailed classification.
@@ 7.3 Relationship of BP/X bulge with the stellar mass
In Fig. 4a, we have plotted the stellar mass distribution of both-sided (green), strong both-sided (purple), one-sided (red), and control sample (blue) in this study. We found that the distribution of control galaxies peaks at lower stellar mass than the galaxies with BP/X bulges. The median of the control sample lies at , whereas the sample with both-sided, strong both-sided, and one-sided BP/X bulges have median at , respectively. Our results are consistent with the previous findings of Erwin & Debattista (2017); Kruk et al. (2019); Marchuk et al. (2022), suggesting massive galaxies most likely host the BP/X bulges.
Further, we measure the BP/X fraction in our sample of edge-on galaxies. Fig. 5 shows the fraction of both-sided (green), strong both-sided (purple), one-sided (red) bulges with the stellar mass of galaxies. The vertical bars in respective curve represent statistical uncertainty in measurement defined as , where being fraction of a bulge class at any stellar mass bin and being total number of galaxies in that bin. Similar to the previous studies, we notice increasing BP/X bulge fraction with increasing stellar mass of galaxies. In the stellar mass range shown here, the fraction of one-sided BP/X bulges is always higher than the strong both-sided bulges at any given stellar mass. Additionally, the fraction of both-sided BP/X bulges, which also includes strong both-sided bulges, is always the highest for any stellar mass. In the shown stellar mass range, we can see an increase in strong both-sided, one-sided and both-sided bulge fraction, respectively, from 0, 0, and 20 to 18, 26, and 42 percent.
@@ 7.4 Dependence of BP/X bulge on stellar surface density
Fig. 6 clearly shows that the fraction of any type of BP/X structure increases with the increasing average surface density of galaxies. The low surface density disks () show nearly constant fraction of both-sided and strong both-sided BP/X bulges. Below this surface density limit, the fraction of strong both-sided structures is less than , while both-sided makes about , and the contribution of one-side structures is about . For high surface density disks (), the fraction of BP/X structure hosting disks rises by about in all three groups. This analysis shows that the high surface density disks are more likely to host BP/X structures as compare to the low surface density disks.
@@ 8 Summary
We visually classified 6653 edge-on galaxies into 1673 () both-sided bulges, and 1112 () one-sided bulges. Among both-sided bulges 504 () are strong both-sided with prominent X-structure, and remaining 1169 () show faint X-structure. When combined together, both-sided and one-sided make (2785 galaxies) of the total sample. Our sample provides the largest catalog of BP/X bulges to date, which includes a category of candidate one-sided BP/X bulges. The catalog of one-sided structures is one of its kind sample, which likely contains samples of ongoing buckling events in edge-on galaxies.
The fraction of one-sided BP/X bulges is always higher than the strong both-sided bulges at any given stellar mass. The fraction of both-sided BP/X bulges is always the highest for any stellar mass. At , the fraction of strong both-sided, one-sided and both-sided bulge is 18, 26, and 42 percent, respectively.
@@
This study provides a classification of BP/X bulges in a large sample of edge-on galaxies, along with their statistical properties and evolutionary trends. We demonstrate the effect of bar formation and evolution on the mass–size and mass–density scaling relations. The analysis uses single-component fitting for size measurements and empirical relations for dust correction and mass estimation, which introduce significant scatter in the scaling relations. Future studies employing detailed 2D decomposition for scale size measurements and SED fitting for stellar mass, star formation rate, dust content, and star formation history will be instrumental in the era of large-scale surveys, e.g., Euclid, LSST, 4MOST, and Roman. With the increase in spatial resolution, it will be possible to distinguish between classical bulge and weak BP/X bulge. We can find the real one-sided BP/X bulge galaxies from the class of one-sided galaxies. Additionally, our sample can be used to train AI models for studying large datasets.
@@ Acknowledgements.
- Díaz-García et al. (2020) Díaz-García, S., Moyano, F. D., Comerón, S., et al. 2020, A&A, 644, A38
- Díaz-García et al. (2016) Díaz-García, S., Salo, H., Laurikainen, E., & Herrera-Endoqui, M. 2016, Astronomy & Astrophysics, 587, A160
@@
- Garland et al. (2024) Garland, I. L., Walmsley, M., Silcock, M. S., et al. 2024, Monthly Notices of the Royal Astronomical Society, 532, 2320
@@
- Holmberg (1958) Holmberg, E. 1958, Meddelanden fran Lunds Astronomiska Observatorium Serie II, 136, 1